OpenStudy (anonymous):

At the local ballpark, the team charges $5 for each ticket and expects to make $1,300 in concessions. The team must pay its players $1,800 and pay all other workers $1,500. Each fan gets a free bat that costs the team $3 per bat. How many tickets must be sold to break even?

OpenStudy (texaschic101):

5t + 1300 = 1800 + 1500 + 3t 5t + 1300 = 3300 + 3t 5t - 3t = 3300 - 1300 2t = 2000 t = 2000/2 t = 1000 check... 5t + 1300 = 1800 + 1500 + 3t 5(1000) + 1300 = 1800 + 1500 + 3(1000) 5000 + 1300 = 3300 + 3000 6300 = 6300 (correct) To break even, the team would have to sell 1000 tickets

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

Than you so much!

OpenStudy (texaschic101):

anytime :)